# cadrille: Multi-modal CAD Reconstruction with Online Reinforcement Learning

### v2

@@ 4.4 RL Fine-tuning
Empirically, we found that hard example mining leads to a faster convergence of RL fine-tuning. Consequently, we only use examples $q$ where the reward $R(\tau)$ averaged over three samples produced by the SFT model is less than 7.5.
@@
Direct Preference Optimization (DPO) rafailov2023dpo learns from pairwise preference data, approximating an implicit reward via a reparameterized Bradley-Terry model.
We construct the training dataset by sampling $K=5$ Python codes $\tau$ for each input $q$ from the SFT model $\pi_{\theta_{r}}$ at temperature $T=1.0$. At each training step for the given sample, we randomly select two outputs. The output with a larger reward $R(\tau)$ is considered to be a preferred prediction $\tau_{w}$, and another is non-preferred $\tau_{l}$. The optimization objective is formulated as:
@@
We combine two recent modifications of the popular GRPO shao2024deepseekmath method: Dr. GRPO liu2025dr.grpo which eliminates the need for a reference model and modifies the objective, and CPPO lin2025cppo which uses samples with the strongest signal. The hybrid approach ensures both computational efficiency and accuracy; hereinafter, it is referred to as Dr. CPPO.
$G$ sequences $\{\tau_{g}\}_{g=1}^{G}$ are sampled from the current policy $\pi_{\theta_{\text{old}}}(\tau\mid q)$ for a given input $q$ with temperature $T=1.0$. For each output $g$, the advantage $A_{g}$ is estimated as $A_{g}={r_{g}-\operatorname{mean}(\{r_{i}\}_{i=1}^{G})}$. $N$ samples with the highest $|A_{g}|$ are used to form a batch $\mathcal{B}$ and perform policy update by maximizing PPO schulman2017ppo objective:
@@ 5 Experiments
DeepCAD wu2021deepcad (denoted as D in Tables) serves as our primary benchmark for supervised training. We adopt the Text2CAD version of DeepCAD, which enriches it with textual descriptions. The training set comprises approximately 160k samples, while 8046 are left for testing.
For SFT, we also use the procedurally generated CAD-Recode rukhovich2024cad-recode dataset (denoted as R). It is an order of magnitude larger than DeepCAD, consisting of approximately 1 million CAD programs written in CadQuery cadquery , a parametric Python-based CAD language.
Fusion360 willis2021fusion360 (denoted as F) is a small CAD reconstruction benchmark with complex and realistic CAD models. In the standard experimental protocol, only the test subset (1725 samples) is used, as the lack of Python CAD sequences makes it unsuitable for conventional supervised training. Still, we can use its training set (6900 samples) for our annotation-free RL fine-tuning.
To show the versatility and applicability of our approach, in addition to handcrafted and procedurally generated meshes, we report metrics on the CC3D dataset mallis2023cc3d . It features 2973 input point clouds sampled from real scans of CAD models with noisy values, missing parts, and smoothed edges.
@@
Following CAD-Recode rukhovich2024cad-recode , we evaluate the quality of the predicted CAD models using three metrics: Chamfer Distance (CD), Intersection over Union (IoU), and Invalidity Ratio (IR). Since invalid CAD models introduce a notable bias into mean estimates, we report both mean and a more robust median CD, both computed using 8192 points. CD values are multiplied by 103. The divergence between ground truth and reconstructed meshes is measured using IoU (in %). The IR indicates the percentage of generated sequences that do not produce a valid CAD model.
@@ 5.1 Supervised Fine-Tuning
In Tab. 1, we compare cadrille with single-modality CAD reconstruction methods on DeepCAD. Here, we specify input modalities with subscripts: p stands for point clouds, i denotes images, and t is for texts. cadrille trained jointly on point clouds, multi-view images, and texts from the DeepCAD training set (Dpit) outperforms the modality-specific baselines. Noticeably, IR is reduced almost twice for point clouds (from 1.1 to 0.4) and 7x for images (3.6 to 0.5).
@@
Both Fusion360 and CC3D datasets do not provide annotations in a compatible format, and are only used for testing in the standard evaluation protocol khan2024cad-signet . Accordingly, testing on these datasets is performed in a zero-shot scenario, which allows assessing the generalization ability of CAD reconstruction approaches. Furthermore, since CC3D contains real scans of objects, this experiment emulates real-world application.
We report quality of image-based and point cloud-based CAD reconstruction in Tab. 2 and  3, respectively. CADCrafter chen2025cadcrafter is the only method performing CAD reconstruction based on multi-view images. However, the authors of CADCrafter only report metrics on the DeepCAD dataset, and benchmarking it on other datasets is problematic since the code has never been released. To establish a baseline in image-based CAD reconstruction, we combine two off-the-shelf state-of-the-art methods, namely, multi-view reconstruction method LRM hong2024lrm ; xu2024instantmesh and CAD-Recode. LRM takes multi-view images as inputs and produces a mesh, which is turned into a point cloud via surface sampling, and this point cloud is then passed to CAD-Recode to create a CAD model. As can be seen in Tab. 2, cadrille  trained on CAD-Recode outperforms both baselines.
In Tab. 3, we compare cadrille  against the state-of-the-art approaches originally trained on the DeepCAD (CAD-SIGNet khan2024cad-signet ) and CAD-Recode rukhovich2024cad-recode datasets. As could be expected, cadrille  is on par with CAD-Recode, while delivering substantially better quality w.r.t. CAD-SIGNet.
@@
PointNet→DeepCAD wu2021deepcad Dp 9.64 46.7 7.1
Point-BERT→HNC-CAD xu2023hnc-cad Dp 8.64 65.3 5.6
MultiCAD ma2023multicad Dp 8.09 11.5
TransCAD dupont2024transcad Dp 4.51 65.5 1.1
PrismCAD lambourne2022prismcad Dp 4.28 72.1 16.2
Point2Cyl uy2022point2cyl Dp 4.27 73.8 3.9
CAD-Diffuser ma2024cad-diffuser Dp 3.02 74.3 1.5
CAD-SIGNet khan2024cad-signet Dp 0.29 77.3 5.0
DINOv2→HNC-CAD xu2023hnc-cad Di 2.08 10.1
DINOv2→DeepCAD wu2021deepcad Di 1.13 10.6
CADCrafter chen2025cadcrafter Di 0.26 3.6
BERT→DeepCAD wu2021deepcad Dt 32.82 10.0
Text2CAD khan2024text2cad Dt 0.37 71.5 0.9
@@
CAD-SIGNet khan2024cad-signet Dp 0.29 77.3 5.0 0.70 58.4 9.3 4.42 39.1 15.5
CAD-Recode rukhovich2024cad-recode Rp 0.18 87.1 3.1 0.19 79.1 5.0 0.54 60.5 9.8
@@
Table 3: Results of CAD reconstruction from point clouds. cadrille performs on par with CAD-Recode rukhovich2024cad-recode when trained on the CAD-Recode dataset (R). With RL, cadrille establishes state-of-the-art on DeepCAD, Fusion360 and real-world CC3D.
@@ 5.2 Reinforcement Learning
Fine-tuning cadrille  using offline DPO reduces IR twice in most cases, while accuracy scores are not affected (rows 3 and 5 in both Tables). In the meantime, Dr. CPPO beats SFT in terms of all metrics, adding 3-9% to IoU scores and bringing IR under 0.2% in all benchmarks (row 6). The observed improvement of CAD reconstruction accuracy aligns well with the experimental results obtained in other tasks where feedback can be programmatically computed shao2024deepseekmath .
@@
6 Limitations
Despite strong performance, our method has several limitations. It relies on either relatively small and limited hand-crafted benchmarks (DeepCAD), or large-scale procedurally generated datasets (CAD-Recode), which may not reflect the complexity of real-world CAD data. Due to the unavailability of CC3D’s training split in the public domain, training or annotation-free fine-tuning on real data remains underexplored. Additionally, the text modality is underutilized, since natural language descriptions are scarce, and VLM-generated captions from Text2CAD often lack realism and might impose risks of data leakage due to being overly descriptive and detailed. Moreover, while cadrille can process all three modalities, it treats them independently and lacks mechanisms to compensate for low-quality or missing inputs. Besides, we only use images for RL fine-tuning. Tuning on point clouds appeared to be unstable, which we attribute to 1) point clouds being a non-native input modality for an LLM and 2) suboptimal point cloud processing strategies inherited from CAD-Recode.
@@ 7 Conclusion
We introduced cadrille, a multimodal CAD reconstruction model that is capable of processing point clouds, multi-view images, and text inputs within a unified VLM-based framework. By adopting a two-stage training paradigm, namely, supervised fine-tuning on synthetic data followed by reinforcement learning fine-tuning with programmatic feedback, we improved both reconstruction quality and validity ratio. Our empirical study demonstrated that online RL approaches are especially beneficial in the CAD reconstruction scenario. The proposed method achieves new state-of-the-art results in multiple benchmarks, including a real-world dataset, highlighting its robustness, generalizability, and potential for further use in applications.
@@ Appendix A Qualitative Results
The pipeline consists of three steps. First, an input image is processed using the recent image-to-mesh InstantMesh xu2024instantmesh method, that produces a mesh. Second, we follow the same protocol as in other experiments to convert this mesh to a point cloud 6 or four multi-view images 7. We claim that the obtained results look promising, and this practical pipeline opens new opportunities of in-the-wild CAD reconstruction.
@@
Text-based CAD reconstruction
Fig. 8 shows results of text-based CAD reconstruction on the DeepCAD dataset. Long textual descriptions still cannot define even simple 3D shapes comprehensively and unambiguously, making CAD reconstruction from textual inputs the most challenging. Both Text2CAD khan2024text2cad and cadrille struggle to recover correct geometry, yet our approach yields more accurate predictions, which is also reflected in the quantitative metrics.
Text Text2CAD cadrille GT
@@ Appendix B Quantitative Results
Data for RL fine-tuning
As described in Sec. 4.4 in the main paper, not all available data is used for RL fine-tuning, but only the most hard examples are considered. By varying the threshold, we can adjust the difficulty level and the ratio of data selected for fine-tuning. Finding the proper balance is crucial for RL fine-tuning, since it has a notably larger time- and memory footprint compared to SFT, and is hardly feasible without hard example mining.
RL Data DeepCAD Fusion360 CC3D
DeepCAD Fusion360 CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓
10k 0.7k 0.18 89.0 1.0 0.18 81.0 2.6 0.74 59.8 6.3
20k 1.4k 0.17 90.4 0.7 0.18 82.3 0.9 0.63 62.4 4.0
30k 2k 0.17 90.9 0.2 0.17 83.5 0.5 0.62 62.3 0.9
40k 2.5k 0.17 91.2 0.1 0.17 83.9 0.2 0.60 63.3 0.7
50k 3k 0.17 92.2 0.0 0.17 84.6 0.0 0.57 65.0 0.1
@@
Mean CD is sometimes reported alongside median CD. In the main paper, we only provided median CD as the most common CAD reconstruction metric. Nevertheless, the evaluation protocol might be considered incomplete without mean CD. In Tab. 5, we report mean CD metric for all scenarios on all three datasets. Our SFT model outperforms all competitors. The gain is especially tangible in text-based CAD reconstruction, where Text2CAD mean CD is reduced 6x from 26.4 to 3.95. With RL fine-tuning (Dr. CPPO) the new state-of-the-art is set in both image-based and point cloud-based CAD reconstruction on all three datasets. The most dramatic improvement is demonstrated on the Fusion360 dataset, where the relative increase exceeds 40% (Tab. 6).
@@
PointNet→DeepCAD wu2021deepcad Dp 42.5 330
CAD-SIGNet khan2024cad-signet Dp 6.81 14.5 32.6
CAD-Recode rukhovich2024cad-recode Rp 0.83 1.21 3.21
@@
BERT→DeepCAD wu2021deepcad Dt 97.9
Text2CAD khan2024text2cad Dt 26.4
@@
More baselines on Fusion360
In Tab. 6, we compare cadrille  to more point cloud-based CAD reconstruction baselines on the Fusion360 dataset. As can be observed, cadrille  outperforms competitors trained either on DeepCAD or CAD-Recode datasets.
Inference time
When target at practical use, efficiency is an issue. In Tab. 7, we compare inference time of our multi-modal cadrille  against previous best single-modal methods: CAD-Recode rukhovich2024cad-recode , Text2CAD khan2024text2cad , and our baseline LRM→CAD-Recode.
Our cadrille  is built on top of Qwen2-VL-2B, the smallest Qwen2 model with vision capabilities. When inferred on point clouds, it is 20% slower compared to CAD-Recode using Qwen2-1.5B. The image inference takes comparable time to proceed. Processing text prompts from Text2CAD dataset lasts notably longer, so that the inference time almost doubles and reaches 3.9 seconds. Text2CAD uses a smaller and faster BERT-large model, that allows achieving efficiency at cost of accuracy. Compared to Text2CAD, cadrille is delivers 6x better mean CD 5 being only 2x slower.
Method Train CD↓ IoU↑ IR↓
DeepCAD wu2021deepcad Dp 89.2 39.9 25.2
MultiCAD ma2023multicad Dp 42.2 16.5
HNC-CAD xu2023hnc-cad Dp 36.8 63.5 7.3
TransCAD dupont2024transcad Dp 33.4 60.2 2.4
CAD-Diffuser ma2024cad-diffuser Dp 3.85 63.2 1.7
CAD-SIGNet khan2024cad-signet Dp 0.70 58.3 9.3
cadrille Dp 0.66 63.7 0.6
CAD-Recode rukhovich2024cad-recode Rp 0.19 79.1 5.0
cadrille Rp 0.19 79.8 2.8
Method LLM Points Images Text
CAD-Recode rukhovich2024cad-recode Qwen2-1.5B 1.8
LRM→CAD-Recode Qwen2-1.5B 4.4
Text2CAD khan2024text2cad BERT-336M 1.7
cadrille Qwen2-VL-2B 2.0 2.0 3.9
Single-image CAD reconstruction
To compare against single-image CAD reconstruction methods, namely, CAD-GPT wang2025cad-gpt and Img2CAD chen2024img2cad , we conduct an experiment with single-view images on the DeepCAD test set. Our SFT model improves over CADCrafter chen2025cadcrafter , reducing median CD from 0.72 to 0.21. As could be expected, the results of single-view CAD reconstruction are slightly inferior to the ones obtained from multi-view images (81% vs 86% IoU).
@@
GPT-4o hurst2024gpt-4o ; wang2025cad-gpt 62.6 64.4
CAD-GPT wang2025cad-gpt 9.77 1.6
DINOv2→HNC-CAD xu2023hnc-cad 2.14 11.4
Img2CAD chen2024img2cad 1.60 28.8
DINOv2→DeepCAD wu2021deepcad 1.26 12.3
CADCrafter chen2025cadcrafter 0.72 8.1
@@
Image-based CAD reconstruction can be performed in zero-shot way using existing VLMs. CAD-GPT wang2025cad-gpt sets a weak baseline using GPT-4o, that has an invalidity ratio of 64% Tab. 8
We construct another baseline for CAD reconstruction from multi-view images. Four images rendered with orthogonal viewing directions are given to GPT-o4-mini to produce a Python code of CAD model. We apply iterative closest point (ICP) to align predictions before computing metrics so that correct predictions with wrong orientation are not penalized. As can be seen in Tab.  9, this strategy allows achieving significantly better results compared to CAD-GPT. Still, invalidity ratio is as high as 15% and IoU is 30% lower compared to our cadrille.
Appendix C Implementation Details

### v3

@@ 4.4 RL Fine-tuning
Empirically, we found that hard example mining leads to a faster convergence of RL fine-tuning. Consequently, we only use examples $q$ where the reward $R(\tau)$ averaged over three samples produced by the SFT model is less than $R_{\text{th}}$, where $R_{\text{th}}=7.5$.
@@
Direct Preference Optimization (DPO) Rafailov et al. (2023) learns from pairwise preference data, approximating an implicit reward via a reparameterized Bradley-Terry model.
We construct the training dataset by sampling $K=5$ Python codes $\tau$ for each input $q$ from the SFT model $\pi_{\theta_{r}}$. At each training step for the given sample, we randomly select two outputs. The output with a larger reward $R(\tau)$ is considered to be a preferred prediction $\tau_{w}$, and another is non-preferred $\tau_{l}$. The optimization objective is formulated as:
@@
We combine two recent modifications of the GRPO: Dr. GRPO Liu et al. (2025b) which eliminates the need for a reference model and modifies the objective, and CPPO Lin et al. (2025) which uses samples with the strongest signal. The hybrid approach ensures both computational efficiency and accuracy; hereinafter, it is referred to as Dr. CPPO.
$G$ sequences $\{\tau_{g}\}_{g=1}^{G}$ are sampled from the current policy $\pi_{\theta_{\text{old}}}(\tau\mid q)$ for a given input $q$ with temperature $T=1.0$. For each output $g$, the advantage $A_{g}$ is estimated as $A_{g}={r_{g}-\operatorname{mean}(\{r_{i}\}_{i=1}^{G})}$. $N$ samples with the highest $|A_{g}|$ are used to form a batch $\mathcal{B}$ and perform policy update by maximizing PPO Schulman et al. (2017) objective:
@@ 5 Experiments
DeepCAD (Wu et al., 2021) (denoted as D in Tables) serves as our primary benchmark for supervised training. We adopt the Text2CAD version of DeepCAD, which enriches it with textual descriptions. The train set comprises approximately 160k samples, while 8046 are left for testing.
For SFT, we also use the procedurally generated CAD-Recode (Rukhovich et al., 2024) dataset (denoted as R in the tables). It is an order of magnitude larger than DeepCAD, consisting of approximately 1 million CAD programs written in CadQuery (Authors, 2024), a parametric Python-based CAD language.
Fusion360 (Willis et al., 2021) (denoted as F) is a small CAD reconstruction benchmark with complex and realistic CAD models. In the standard evaluation protocol, only the test split (1725 samples) is used, as absence of Python CAD sequences makes it unsuitable for conventional supervised training. Still, we can use its train set (6900 samples) for our annotation-free RL fine-tuning.
To show the versatility and applicability of our approach, in addition to handcrafted and procedurally generated meshes, we report metrics on the real-world CC3D dataset (Mallis et al., 2023) (denoted as C). It contains 2973 point clouds sampled from scans of CAD models with noisy values, missing parts, and smoothed edges.
Omni-CAD (denoted as O) presented in the CAD-MLLM paper (Xu et al., 2024b) is a large-scale dataset of handcrafted CAD models, sourced from the web. We evaluate on its test split composed of over 27K samples, and report results in Appendix B.
@@
Following CAD-Recode, we normalize ground truth CAD models so that they fit into $0.5,0.5]^{3}$ and report three metrics: Chamfer Distance (CD), Intersection over Union (IoU), and Invalidity Ratio (IR). Since invalid CAD models introduce a notable bias into mean estimates, we report median CD computed using 8192 points. CD values are multiplied by 103. The divergence between ground truth and reconstructed meshes is measured using IoU (in %). The IR indicates the percentage of generated sequences that do not produce a valid CAD model.
@@ 5.1 Supervised Fine-Tuning
In Tab. 1, we compare cadrille with single-modal CAD reconstruction methods on DeepCAD. Here, input modalities are denoted with subscripts: p stands for point clouds, i for images, and t for texts. cadrille trained jointly on point clouds, multi-view images, and texts from the DeepCAD training set (Dpit) outperforms the modality-specific baselines. Noticeably, IR is reduced almost twice for point clouds (from 1.1 to 0.4) and 7x for images (3.6 to 0.5).
@@
Both Fusion360 and CC3D datasets do not provide annotations in a compatible format, and are only used for testing in the standard evaluation protocol (Khan et al., 2024a). Accordingly, testing on these datasets is performed in a zero-shot scenario, which allows assessing the generalization ability of CAD reconstruction approaches. Furthermore, since CC3D contains real scans of objects, this experiment emulates real-world application.
We report CAD reconstruction quality from images and point clouds in Tab. 2 and  3, respectively. CADCrafter is the only method performing CAD reconstruction based on multi-view images. However, the authors of CADCrafter only report metrics on the DeepCAD dataset, and benchmarking it on other datasets is problematic since the code has not been released. To establish a baseline in image-based CAD reconstruction, we combine two off-the-shelf state-of-the-art methods, namely, multi-view reconstruction method LRM Hong et al. (2024); Xu et al. (2024a) and CAD-Recode. LRM takes multi-view images as inputs and produces a mesh, which is turned into a point cloud via surface sampling, and this point cloud is then passed to CAD-Recode to create a CAD model. As can be seen in Tab. 2, cadrille  trained on CAD-Recode outperforms both baselines.
In Tab. 3, we compare cadrille  against the state-of-the-art approaches originally trained on the DeepCAD (CAD-SIGNet) and CAD-Recode datasets. As could be expected, cadrille  is on par with CAD-Recode, while delivering substantially better quality w.r.t. CAD-SIGNet.
@@
PointNet→DeepCAD Dp 9.64 46.7 7.1
Point-BERT→HNC-CAD Dp 8.64 65.3 5.6
MultiCAD Dp 8.09 11.5
TransCAD Dp 4.51 65.5 1.1
PrismCAD Dp 4.28 72.1 16.2
Point2Cyl Dp 4.27 73.8 3.9
CAD-Diffuser Dp 3.02 74.3 1.5
CAD-SIGNet Dp 0.29 77.3 5.0
DINOv2→HNC-CAD Di 2.08 10.1
DINOv2→DeepCAD Di 1.13 10.6
CADCrafter Di 0.26 3.6
BERT→DeepCAD Dt 32.82 10.0
CADmium Dt 0.38 4.3
Text2CAD Dt 0.37 71.5 3.7
CAD-Coder Dt 0.33 5.3
Text-to-CadQuery Dt 0.22 1.3
@@
CAD-SIGNet Dp 0.29 77.3 5.0 0.70 58.4 9.3 4.42 39.1 15.5
CAD-Recode Rp 0.18 87.1 3.1 0.19 79.1 5.0 0.54 60.5 9.8
@@
Table 3: Results of CAD reconstruction from point clouds. cadrille performs on par with CAD-Recode when trained on the CAD-Recode dataset (R). With RL, cadrille establishes state-of-the-art on DeepCAD, Fusion360 and real-world CC3D.
@@ 5.2 Reinforcement Learning
Fine-tuning cadrille  using offline DPO reduces IR twice in most cases, while accuracy scores are not affected (rows 3 and 5 in both Tables). In the meantime, Dr. CPPO beats SFT in terms of all metrics, adding 3-9% to IoU scores and bringing IR under 0.2% on all benchmarks (row 6). The observed improvement of CAD reconstruction accuracy aligns well with the experimental results obtained in other tasks where feedback can be programmatically computed Shao et al. (2024).
@@ 6 Conclusion
We introduced cadrille, a multimodal CAD reconstruction model that is capable of processing point clouds, multi-view images, and text inputs within a unified VLM-based framework. By adopting a two-stage training paradigm, namely, supervised fine-tuning on synthetic data followed by reinforcement learning fine-tuning with programmatic feedback, we improved both reconstruction quality and validity ratio. Our empirical study demonstrated that online RL approaches are especially beneficial in the CAD reconstruction scenario. cadrille achieves new state-of-the-art results in 10 CAD reconstruction benchmarks, including a real-world dataset, highlighting its robustness, generalizability, and potential for further use in applications. Based on our study, we identify the following promising research directions for the future work: 1) combine modalities in one prompt to compensate for low-quality or missing inputs 2) perform RL fine-tuning on point clouds, and 3) increase complexity of procedurally generated data and volume of RL fine-tuning data to better adapt to real-world scans.
This work was supported by the The Ministry of Economic Development of the Russian Federation in accordance with the subsidy agreement (agreement identifier  [^00] C313925P4H0002; grant No 139-15-2025-012).
@@ Appendix A Qualitative Results
Text-based CAD reconstruction
Fig. 6 shows results of text-based CAD reconstruction on the DeepCAD dataset. Long textual descriptions still cannot define even simple 3D shapes comprehensively and unambiguously, making CAD reconstruction from textual inputs the most challenging. Both Text2CAD and cadrille struggle to recover correct geometry, yet our approach yields more accurate predictions, which is also reflected in the quantitative metrics.
@@
The pipeline consists of three steps. First, an input image is processed using the recent image-to-mesh InstantMesh Xu et al. (2024a) method, that produces a mesh. Second, we follow the same protocol as in other experiments to convert this mesh to a point cloud 7 or four multi-view images 8. We claim that the obtained results look promising, and this practical pipeline opens new opportunities of in-the-wild CAD reconstruction.
Failure cases
Text Text2CAD cadrille GT
@@ Appendix B Quantitative Results
Results on Omni-CAD
Omni-CAD is claimed to be the first multimodal dataset featuring point clouds, multi-view images, and textual descriptions of CAD models. However, texts are not fully utilized for CAD reconstruction, since generated models are only assessed in a user study and no standard quantitative metrics are provided (Xu et al., 2024b). Accordingly, we validate cadrille only in point cloud-based and image-based scenarios, and report the evaluation results in Tab. 4.
Method RL Point Cloud Multi-view Images
CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓
DeepCAD 45.1 5.7
CAD-MLLM 18.5 1.3 37.7
cadrille 1.00 79.1 3.7 1.22 77.0 6.2
cadrille 0.77 84.2 0.6 0.60 84.8 0.0
Table 4: Results of CAD reconstruction from point clouds and multi-view images from the Omni-CAD dataset. We specify mean CD since it is the only CD metric reported by CAD-MLLM.
Method Train CD↓ IoU↑ IR↓
DeepCAD Dp 89.2 39.9 25.2
MultiCAD Dp 42.2 16.5
HNC-CAD Dp 36.8 63.5 7.3
TransCAD Dp 33.4 60.2 2.4
CAD-Diffuser Dp 3.85 63.2 1.7
CAD-SIGNet Dp 0.70 58.3 9.3
cadrille Dp 0.66 63.7 0.6
CAD-Recode Rp 0.19 79.1 5.0
cadrille Rp 0.19 79.8 2.8
The metrics of prior works are cited in accordance to CAD-MLLM; we multiply their values by 10 for consistency, since they report all scores at 102 scale, while we use 103 scale in all tables. Here, CAD-MLLM results are obtained through training on the Omni-CAD dataset, while cadrille  is evaluated in a pure cross-dataset scenario without any adaptation. In general, the results on Omni-CAD are similar to the ones obtained on other benchmarks, namely, cadrille achieves the highest IoU score among all tested approaches, while IR is the lowest with the large margin, reaching 0.0 on image-based benchmark due to RL fine-tuning.
More baselines on Fusion360
In Tab. 5, we compare cadrille  to more point cloud-based CAD reconstruction baselines on the Fusion360 dataset. As can be observed, cadrille  outperforms competitors trained either on DeepCAD or CAD-Recode datasets. The only existing three-modal approach CAD-MLLM reports only mean CD for the Fusion360 dataset, so we list it among other approaches in another Tab. 6. As can be observed, CAD-MLLM far behind our method, with mean CD value being two orders of magnitude larger than of cadrille.
@@
Mean CD is sometimes reported alongside median CD. In the main paper, we only provided median CD as the most common CAD reconstruction metric. Nevertheless, the evaluation protocol might be considered incomplete without mean CD. In Tab. 6, we report mean CD metric for all scenarios on all three datasets. Due to randomized point sampling, mean CD is not 0 even for perfectly accurate predictions: i.e., evaluation of ground truth CAD models from the DeepCAD dataset gives $0.16$, Fusion360 – $0.13$ and CC3D – $0.15$; as can be observed, while cadrille  delivers the lowest CD scores, there is room for improvement. Our SFT model outperforms all competitors. The gain is especially tangible in text-based CAD reconstruction, where Text2CAD mean CD is reduced 6x from 26.4 to 3.95. With RL fine-tuning (Dr. CPPO) the new state-of-the-art is set in both image-based and point cloud-based CAD reconstruction on all three datasets. The most dramatic improvement is demonstrated on the Fusion360 dataset, where the relative increase exceeds 40% (Tab. 5).
@@
PointNet→DeepCAD Dp 42.5 76.1
CAD-MLLM Opit 33.9
CAD-SIGNet Dp 6.81 14.5 32.6
CAD-Recode Rp 0.83 1.21 3.21
@@
BERT→DeepCAD Dt 97.9
CAD-Coder Dt 74.6
Text2CAD Dt 26.4
Text-to-CadQuery Dt 11.8
@@
GPT-4o 62.6 64.4
CAD-GPT 9.77 1.6
DINOv2→HNC-CAD 2.14 11.4
Img2CAD 1.60 28.8
DINOv2→DeepCAD 1.26 12.3
CADCrafter 0.72 8.1
@@
Single-image CAD reconstruction
To compare against single-image CAD reconstruction methods, namely, CAD-GPT (Wang et al., 2025b) and Img2CAD (Chen et al., 2024), we conduct an experiment with single-view images on the DeepCAD test set. Our SFT model improves over CADCrafter (Chen et al., 2025), reducing median CD from 0.72 to 0.21. As could be expected, the results of single-view CAD reconstruction are slightly inferior to the ones obtained from multi-view images (81% vs 86% IoU).
@@
Image-based CAD reconstruction can be performed in zero-shot way using existing VLMs. CAD-GPT sets a weak baseline using GPT-4o, that has an invalidity ratio of 64% Tab. 7
We construct another baseline for CAD reconstruction from multi-view images. Four images rendered with orthogonal viewing directions are given to GPT-o4-mini to produce a Python code of CAD model. We apply iterative closest point (ICP) to align predictions before computing metrics so that correct predictions with wrong orientation are not penalized. As can be seen in Tab.  8, this strategy allows achieving significantly better results compared to CAD-GPT. Still, invalidity ratio is as high as 15% and IoU is 30% lower compared to our cadrille.
Inference time
When target at practical use, efficiency is an issue. In Tab. 9, we compare inference time of our multimodal cadrille  against previous best single-modal methods: CAD-Recode, Text2CAD, and our baseline LRM→CAD-Recode.
Our cadrille  is built on top of Qwen2-VL-2B, the smallest Qwen2 model with vision capabilities. When inferred on point clouds, it is 20% slower than CAD-Recode using Qwen2-1.5B. The image inference takes comparable time to proceed. Processing text prompts from Text2CAD dataset lasts notably longer, so that the inference time almost doubles and reaches 3.9 seconds. Text2CAD uses a smaller and faster BERT-large model, that allows achieving efficiency at cost of accuracy. Compared to Text2CAD, cadrille delivers 6x better mean CD 6 being only 2x slower.
Appendix C Ablation Experiments
RL fine-tuning vs test-time sampling?
A natural concern about the RL fine-tuning is that it might compromise the diversity of responses, hence affecting the final prediction quality. To address this, we conduct an ablation study to investigate how the number of test-time samples affects the performance of both CAD-Recode and cadrille. As reported in 11, CAD-Recode produces more accurate results with an increasing number of samples, but remains inferior to cadrille. On all the benchmarks, cadrille with 1 sample consistently outperforms CAD-Recode with 2 samples in both accuracy and invalidity ratio. In terms of IR, cadrille with 1 sample is better than CAD-Recode with as many as 10 samples, which actually proves RL to boost the robustness of the model.
As could be expected, the gap between IoU@1 with and without RL being significantly larger than the gap between IoU@k (e.g. $k=10$) with and without RL. This result aligns perfectly with the recent evidence coming from the math domain (see Fig. 2 of Yue et al. (2025) and Fig. 4 of Liu et al. (2025a)).
Method LLM Points Images Text
CAD-Recode Qwen2-1.5B 1.8
LRM→CAD-Recode Qwen2-1.5B 4.4
Text2CAD BERT-336M 1.7
cadrille Qwen2-VL-2B 2.0 2.0 3.9
$K$ DeepCAD Fusion360 CC3D
2 77.3 70.0 49.7
3 86.2 77.8 55.6
5 86.9 78.5 56.0
Table 10: Results of image-based CAD reconstruction on three datasets, obtained with varying number of samples $K$ in DPO. IoU scores are reported.
DeepCAD Fusion360 CC3D
# samples 1 2 10 1 2 10 1 2 10
IoU↑ IR↓ IoU↑ IR↓ IoU↑ IR↓ IoU↑ IR↓ IoU↑ IR↓ IoU↑ IR↓ IoU↑ IR IoU↑ IR↓ IoU↑ IR↓
CAD-Recode 87.1 3.1 89.9 0.9 92.0 0.4 79.1 5.0 83.6 1.3 87.8 0.5 60.5 9.8 64.5 1.7 74.2 0.3
cadrille 90.2 0.0 91.6 0.0 93.1 0.0 85.0 0.2 86.8 0.1 89.1 0.0 67.9 0.2 70.9 0.2 74.7 0.1
Table 11: Results of point-based CAD reconstruction with test-time sampling.
Data for RL fine-tuning
As described in Sec. 4.4 in the main paper, not all available data is used for RL fine-tuning, but only the most hard examples are considered. By varying the hard-mining threshold $R_{\text{th}}$, we control the difficulty level and the ratio of data selected for fine-tuning. Finding the proper balance is crucial for RL fine-tuning, since it has a notably larger time- and memory footprint compared to SFT, and is hardly feasible without hard example mining.
$R_{\text{th}}$ RL Data DeepCAD Fusion360 CC3D
DeepCAD Fusion360 CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓
1 10k 0.7k 0.18 89.0 1.0 0.18 81.0 2.6 0.74 59.8 6.3
3 20k 1.4k 0.17 90.4 0.7 0.18 82.3 0.9 0.63 62.4 4.0
5 30k 2k 0.17 90.9 0.2 0.17 83.5 0.5 0.62 62.3 0.9
6.2 40k 2.5k 0.17 91.2 0.1 0.17 83.9 0.2 0.60 63.3 0.7
7.5 50k 3k 0.17 92.2 0.0 0.17 84.6 0.0 0.57 65.0 0.1
Number of samples in DPO
We investigate how the number of samples used for RL fine-tuning affects the final results. In Tab. 10, we report IoU scores for image-based reconstruction on all three datasets after 20 epochs of DPO fine-tuning. Performance is unstable with only $K=2$ samples, while our default value of $K=5$ yields the best performance. The difference between $K=3$ and $K=5$ is less than 1% of IoU, suggesting that the model saturates with few samples, so 5 samples are sufficient.
Appendix D Implementation Details